The ratio of the width to length of a rectangular patio is 2:3. If the length of the patio is 6 yards, find the area of the patio in square feet.
Answer by TimothyLamb(4379)

---
y/x = 2/3
x = 6*3 ft
---
put the system of linear equations into standard form:
---
y = (2/3)x
x = 18
---
(2/3)x - y = 0
x = 18
---
0.666666x - y = 0
x = 18
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = 18 ft
y = 12 ft
---
area = xy = 18*12
area = xy = 216 sq.ft
---
now by substitution:
---
y = (2/3)x
x = 18
---
y = (2/3)*18
y = 2*6
y = 12
---
solution by substitution method:
x = 18 ft
y = 12 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php